OpenStudy (anonymous):

what is 2110 div. by 10 times the product of 226 div.9 ? ( can be remandirs and simifly completly.

OpenStudy (anonymous):

please help.

OpenStudy (anonymous):

\[\frac{2110}{10 \times \frac{226}{9}} = \frac{2110}{\frac{2260}{9}} = \frac{2110}{1} \times \frac{9}{2260} = \frac{18990}{2260}\]

OpenStudy (anonymous):

thank you.